# On inverse problem for generalized Kerr-type nonlinearities

### v1

@@ Abstract.
We study the inverse problem of reconstructing the shape of unknown inclusions in semilinear elliptic equations with nonanalytic nonlinearities, by extending Ikehata’s enclosure method to accommodate such nonlinear effects. To address the analytical challenges, we construct an approximate solution based on the linearized equation, enabling the enclosure method to operate in this setting. The proposed method applies to a broad class of semilinear elliptic equations with non-analytic nonlinearities, including representative examples such as the Kerr-type nonlinearity, which appears in models of nonlinear optics, and the Ginzburg-Landau-type nonlinearity, which models light propagation in nonlinear dissipative media.
@@ 1. Introduction
Here, the nonlinear term $q(x,\lvert u\rvert)$ in our model is non-analytic in $u$, and is designed to represent a range of physically relevant responses. Precisely, we consider the following structural setting. Assume there exist $0\geq q_{0}\in L^{\infty}(\Omega)$, $q_{1}\in L^{\infty}(\Omega)$, a constant $C_{*}>0$, an integer $\alpha_{1}\geq 2$, a real number $\alpha_{2}>\alpha_{1}$ such that the nonlinear term $q$ satisfies
@@
which arises in models of nonlinear optics, see e.g. [Boy20, MN19] for a physical description. Moreover, this equation arises as the continuum limit of a family of discrete nonlinear Schrödinger equations with long-range lattice interactions, modeling charge transport in biopolymers such as DNA, see e.g. [KLS13]. Using Lemma 2.2 below, one can readily verify that another example of (1.1b) is the Ginzburg-Landau-type nonlinearity, corresponding to the special case
@@
which arises in models of light propagation in nonlinear dissipative media, see e.g. [QMM+20].
@@
The method relies on two main components: the special solutions and the indicator functional. In many classical cases, the special solutions are taken to be complex geometrical optics (CGO) solutions. However, in the previous work [Kua24] involving nonlinear equations with power-type nonlinearities, the CGO solutions were replaced by so-called approximate solutions, which were constructed via a Taylor expansion of the associated solution operator. This construction was inspired by the higher-order linearization framework proposed in [LLLS21], and adapted to the setting of the enclosure method.
A main difficulty in our setting is the lack of analyticity in the nonlinear term $z\mapsto q(x,\lvert z\rvert)z$, which makes the Taylor-based construction of approximate solutions no longer applicable. To address this, we introduce a new strategy for constructing approximate solutions suitable for our case. Although this construction differs conceptually from [Kua24], similar types of estimates are still needed to justify the method. These rely on key inequalities developed in [GKM22, KW23], which play a critical role in establishing the required bounds. Further details of this approach are provided in Section 3.
@@
To formulate the inverse problem in terms of boundary data, it is essential to ensure that the forward problem is well-posed. In particular, the solution to the boundary value problem must exist uniquely and depend continuously on the boundary data, so that the associated Dirichlet-to-Neumann (DtN) map is well-defined. A detailed discussion of the well-posedness is provided in Section 2.
@@
see e.g. [KP88, Kua24] or [Maz11, Section 1.8]. Using Lemma 2.3 below, one can identify a small parameter
@@
Theorem 1.1 (see Theorem 3.6 below for a detailed statement).
@@
It is worth noting that our method does not require the exact expression of the nonlinearity $q$, and in fact, we do not even assume knowledge of the background coefficients associated with the higher-order nonlinear terms (i.e., the $\alpha_{2}$ term). This allows for greater flexibility in the admissible forms of $q$.
@@
In Section 2, we establish the well-posedness of the boundary value problem (1.1), followed by the proof of our main result (Theorem 1.1) in Section 3.
@@ 2. Forward problem
Before tackling the inverse problem, we first consider the forward problem by establishing the well-posedness of the boundary value problem (1.1). We now introduce the following assumption, see also [GKM22, Assumption 2.1]:
@@
The following lemma, which demonstrates that 2.1 generalizes (1.1b) (with $\alpha=\alpha_{1}$), plays a central role in our proof.
@@
(2.2) $-\Delta v_{*}=0,\quad\left.v_{*}\right|_{\partial\Omega}=\lVert f\rVert_{L^{\infty}(\partial\Omega)},$
@@
We now see that $v_{*}\pm v_{f}\in H_{0}^{1}(\Omega)$ is the unique solution to
$-(\Delta+q_{0})(v_{*}\pm v_{f})=-q_{0}v_{*}\text{ in $\Omega$.}$
Since $q_{0}\leq 0$ a.e. in $\Omega$, then from (2.4) we see that $-q_{0}v_{*}\geq 0$ a.e. in $\Omega$, and again using maximum principle we see that $v_{*}\pm v_{f}\geq 0$ a.e. in $\Omega$, that is,
@@
For each $F\in L^{2}(\Omega)$, again one can use the Lax-Miltram theorem to guarantee that there exists a unique solution $w_{F}\in H_{0}^{1}(\Omega)$ of
@@
$-\Delta w_{*}=\lVert F\rVert_{L^{\infty}(\Omega)}\text{ in $\Omega$},\quad\left.w_{*}\right|_{\partial\Omega}=0.$
@@
$-(\Delta+q_{0})(w_{*}\pm w_{F})=\lVert F\rVert_{L^{\infty}(\Omega)}\mp F-q_{0}w_{*}\text{ in $\Omega$},\quad\left.\left(w_{*}\pm w_{F}\right)\right|_{\partial\Omega}=0.$
@@
We are now ready to prove the well-posedness of (1.1a) under 2.1 following the ideas in [KW23, Theorem 2.1], which is based on the contraction mapping theorem.
@@
Let $\Omega$ be a bounded $C^{1,1}$-domain, let $0\geq q_{0}\in L^{\infty}(\Omega)$ and let $q$ satisfies 2.1. There exists a sufficiently small $\delta_{0}=\delta_{0}(\alpha,C_{\rm Kerr},\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\in(0,1)$ such that the following statement holds true: if $f\in U_{\delta_{0}}$, then there exists a unique solution $u_{f}\in\mathcal{A}^{2}(\Omega)$ to (1.1a) satisfying
@@
By using (2.16), and consequently using 2.1 with the choice $z_{1}=w+v_{f}$ and $z_{2}=0$, we see that
@@
(2.21) $C_{\rm Kerr}C^{1+\alpha}(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{2+\alpha}\delta_{0}^{\alpha}\leq 1,$
@@
Since $\mathcal{S}$ is linear, by using (2.16) and consequently using 2.1 with the choice $z_{1}=w_{2}+v_{f}$ and $z_{2}=w_{1}+v_{f}$ we see that
@@ 3. Proof of the main result
To clearly present our ideas, we divide the proof of Theorem 1.1 into several subsections.
@@ 3.1. Approximate solutions
In this subsection, we construct approximate solutions suitable for the non-analytic setting. Since the nonlinearity is not analytic, the Taylor-based approach in [Kua24] does not apply; instead, we evaluate the nonlinear term at the linearized solution $v_{f}$, the unique solution of (2.1). Letting $\tilde{q}(x,\lvert z\rvert)=q(x,\lvert z\rvert)-q_{0}(x)$ for all $x\in\Omega$ and $z\in\mathbb{C}$, we define $\tilde{u}_{f}$ as the unique solution to
@@
This substitution may be seen as a conceptual iteration that transforms the original nonlinear equation into a linear one.
@@
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfies (1.1b). Then there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
@@
where $\delta_{0}\in(0,1)$ denotes the constant specified in Lemma 2.3.
@@
$(\Delta+q_{0})(u_{f}-\tilde{u}_{f})=-\left(\tilde{q}(x,\lvert u_{f}\rvert)u_{f}-\tilde{q}(x,\lvert v_{f}\rvert)v_{f}\right)\text{ in $\Omega$},\quad\left.(u_{f}-\tilde{u}_{f})\right|_{\partial\Omega}=0,$
using (2.16) and (1.1b) we see that
@@
(2.16)${\leq}C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\left(\begin{aligned} &\lVert\tilde{q}(\cdot,\lvert u_{f}\rvert)u_{f}-\tilde{q}(\cdot,\lvert v_{f}\rvert)v_{f}-q_{1}(\cdot)(\lvert u_{f}\rvert^{\alpha_{1}}u_{f}-\lvert v_{f}\rvert^{\alpha_{1}}v_{f})\rVert\\ &+\lVert q_{1}(\cdot)(\lvert u_{f}\rvert^{\alpha_{1}}u_{f}-\lvert v_{f}\rvert^{\alpha_{1}}v_{f})\rVert_{L^{\infty}(\Omega)}\end{aligned}\right)$ )
Consequently, we use (1.1b) and Lemma 2.2, as well as the stability estimates in (2.17) and (2.18), to see that
@@ 3.2. An indicator functional and its approximation
We define $q_{b}(x,\lvert z\rvert):=q_{0}(x)+q_{1,b}\lvert z\rvert^{\alpha_{1}}$ for all $x\in\Omega$ and $z\in\mathbb{C}$, where $q_{1,b}\in L^{\infty}(\Omega)$ is the function given in (1.5a). By using Lemma 2.3, there exists a $\delta_{0}\in(0,1)$ such that the following statement holds true: if $f\in U_{\delta_{0}}$, then there exists a unique solution $u_{b,f}\in\mathcal{A}^{2}(\Omega)$ to
@@
Using Lemma 3.1, we know that there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1,b}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
@@
where $v_{f}$ is the unique solution to (2.1). Based on the indicator functional in [Kua24] and the structure of our current model, we propose the following variant: The existing enclosure method suggests us to consider the following indicator functionals:
$\mathcal{I}(f):=\int_{\partial\Omega}(\partial_{\nu}u_{f}-\partial_{\nu}\tilde{u}_{b,f})\overline{f}\,\mathrm{d}S\quad\text{for all $f\in U_{\delta_{0}}$}$
and
~$\widetilde{\mathcal{I}}(f):=\int_{\partial\Omega}(\partial_{\nu}\tilde{u}_{f}-\partial_{\nu}\tilde{u}_{b,f})\overline{f}\,\mathrm{d}S\quad\text{for all $f\in U_{\delta_{0}}$.}$
@@
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfies (1.1b). Then there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
@@
where $\delta_{0}\in(0,1)$ is the constant given in Lemma 2.3.
Proof. Using the trace theorem [AH09, Theorem 7.3.11] (which requires the condition $\partial\Omega\in C^{1,1}$), for each $f\in U_{\delta_{0}}$, we compute that
@@
Our lemma follows directly from Lemma 3.1. ∎
@@
~$\widetilde{\mathcal{I}}(f)=-\int_{D}q_{1,D}(x)\lvert v_{f}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{f}\rvert)\lvert v_{f}\rvert^{2}\,\mathrm{d}x\quad\text{for all $f\in U_{\delta_{0}}$,}$
where $v_{f}$ is the unique solution to (2.1) and $\delta_{0}\in(0,1)$ denotes the constant specified in Lemma 2.3.
@@
Proof of Lemma 3.3. We see that the function $z_{f}:=\tilde{u}_{f}-\tilde{u}_{b,f}$ satisfies

### v2

# An inverse problem for semilinear elliptic equations with generalized Kerr-type nonlinearities

@@ Abstract.
We study the inverse problem of reconstructing the shape of unknown inclusions in semilinear elliptic equations with nonanalytic nonlinearities, by extending Ikehata’s enclosure method to accommodate such nonlinear effects. To address the analytical challenges, we construct an approximate solution based on the linearized equation, enabling the enclosure method to operate in this setting. Under suitable structural conditions on the nonlinearity, we establish a shape reconstruction result. The proposed method applies to a broad class of semilinear elliptic equations with non-analytic nonlinearities, including representative examples such as the Kerr-type nonlinearity, which appears in models of nonlinear optics, and the Ginzburg-Landau-type nonlinearity, which models light propagation in nonlinear dissipative media.
@@
Acknowledgments
We would like to thank the anonymous readers for their comments and for pointing out several typos in the first draft on arXiv.
@@ 1. Introduction
[…]
@@
which arises in models of nonlinear optics, see e.g. [Boy20, MN19] for a physical description. Moreover, this equation arises as the continuum limit of a family of discrete nonlinear Schrödinger equations with long-range lattice interactions, modeling charge transport in biopolymers such as DNA, see e.g. [KLS13]. Using Lemma˜2.2 below, one can readily verify that another example of (1.1b) is the Ginzburg-Landau-type nonlinearity, corresponding to the special case
@@
which arises in models of light propagation in nonlinear dissipative media, see e.g. [QMM+20]. Furthermore, the nonlinearity considered in [LLST22] is also an instance of the form (1.1b).
@@
The method relies on two main components: the special solutions and the indicator functional. In many classical cases, the special solutions are taken to be complex geometrical optics (CGO) solutions. However, in the previous work [Kua24] involving nonlinear equations with power-type nonlinearities, the CGO solutions were replaced by so-called approximate solutions, which were constructed via a Taylor expansion of the associated solution operator. This construction was inspired by the higher-order linearization framework proposed in [LLLS21a, LLLS21b, KU20], and adapted to the setting of the enclosure method.
A main difficulty in our setting is the lack of analyticity (cf. [LLLS21a, LLLS21b, KU20]) of the nonlinear term $z\mapsto q(x,\lvert z\rvert)z$, which renders the Taylor-based construction of approximate solutions in [Kua24] inapplicable. To address this, we introduce a new strategy for constructing approximate solutions suitable for our case. Although this construction differs conceptually from [Kua24], similar types of estimates are still needed to justify the method. These rely on key inequalities developed in [GKM22, KW23], which play a critical role in establishing the required bounds. Further details of this approach are provided in Section˜3.
@@
To formulate the inverse problem in terms of boundary data, it is essential to ensure that the forward problem is well-posed. In particular, the solution to the boundary value problem must exist uniquely and depend continuously on the boundary data, so that the associated Dirichlet-to-Neumann (DtN) map is well-defined. A detailed discussion of the well-posedness is provided in Section˜2.
@@
see e.g. [KP88, Kua24] or [Maz11, Section 1.8]. Using Lemma˜2.3 below, one can identify a small parameter
@@
Theorem 1.1 (see Theorem˜3.6 below for a detailed statement).
@@
It should be noted that our method does not require the exact expression of the nonlinearity $q$. This allows for greater flexibility in the admissible forms of $q$.
@@
In Section˜2, we establish the well-posedness of the boundary value problem (1.1), followed by the proof of our main result (Theorem˜1.1) in Section˜3.
@@ 2. Forward problem
Before we tackle the inverse problem, we first consider the forward problem by establishing the well-posedness of the boundary value problem (1.1). We now introduce the following assumption, see also [GKM22, Assumption 2.1]:
@@
The following lemma, which demonstrates that ˜2.1 generalizes (1.1b) (with $\alpha=\alpha_{1}$), plays a central role in our proof.
@@
(2.2) $\Delta v_{*}=0,\quad\left.v_{*}\right|_{\partial\Omega}=\lVert f\rVert_{L^{\infty}(\partial\Omega)},$
@@
We now see that $v_{*}\pm v_{f}\in H^{1}(\Omega)$ solves
$(\Delta+q_{0})(v_{*}\pm v_{f})=-q_{0}v_{*}\text{ in $\Omega$.}$
Since $(v_{*}\pm v_{f})|_{\partial\Omega}=\lVert f\rVert_{L^{\infty}(\partial\Omega)}\pm f\geq 0$ and $q_{0}\leq 0$ a.e. in $\Omega$, then from (2.4) we see that $q_{0}v_{*}\geq 0$ a.e. in $\Omega$, and again using maximum principle we see that $v_{*}\pm v_{f}\geq 0$ a.e. in $\Omega$, that is,
@@
For each $F\in L^{2}(\Omega)$, again one can use the Lax-Milgram theorem to guarantee that there exists a unique solution $w_{F}\in H_{0}^{1}(\Omega)$ of
@@
$\Delta w_{*}=\lVert F\rVert_{L^{\infty}(\Omega)}\text{ in $\Omega$},\quad\left.w_{*}\right|_{\partial\Omega}=0.$
@@
$(\Delta+q_{0})(w_{*}\pm w_{F})=\lVert F\rVert_{L^{\infty}(\Omega)}\mp F-q_{0}w_{*}\text{ in $\Omega$},\quad\left.\left(w_{*}\pm w_{F}\right)\right|_{\partial\Omega}=0.$
@@
We are now ready to prove the well-posedness of (1.1a) under ˜2.1 following the ideas in [KW23, Theorem 2.1], which is based on the contraction mapping theorem.
@@
Let $\Omega$ be a bounded $C^{1,1}$-domain, let $0\geq q_{0}\in L^{\infty}(\Omega)$ and let $q$ satisfy ˜2.1. There exists a sufficiently small $\delta_{0}=\delta_{0}(\alpha,C_{\rm Kerr},\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\in(0,1)$ such that the following statement holds true: if $f\in U_{\delta_{0}}$, then there exists a unique solution $u_{f}\in\mathcal{A}^{2}(\Omega)$ to (1.1a) satisfying
@@
By using (2.16), and consequently using ˜2.1 with the choice $z_{1}=w+v_{f}$ and $z_{2}=0$, we see that
@@
(2.21) $C_{\rm Kerr}C^{2+\alpha}2^{1+\alpha}(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})^{2+\alpha}\delta_{0}^{\alpha}\leq 1,$
@@
Since $\mathcal{S}$ is linear, by using (2.16) and consequently using ˜2.1 with the choice $z_{1}=w_{2}+v_{f}$ and $z_{2}=w_{1}+v_{f}$ we see that
@@ 3. Proof of the main result
To clearly present our ideas, we divide the proof of Theorem˜1.1 into several subsections.
@@ 3.1. Approximate solutions
In this subsection, we construct approximate solutions suitable for the non-analytic setting. Since the nonlinearity is not analytic, the Taylor-based approach in [Kua24] does not apply. Instead, our construction adopts an iterative perspective: we take the linearized solution $v_{f}$ as the initial approximation and substitute it directly into the nonlinear term, where $v_{f}$ denotes the unique solution of (2.1).
Precisely, letting $\tilde{q}(x,\lvert z\rvert)=q(x,\lvert z\rvert)-q_{0}(x)$ for all $x\in\Omega$ and $z\in\mathbb{C}$, we define $\tilde{u}_{f}$ as the unique solution to the following linear equation:
@@
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfy (1.1b). Then there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
@@
where $\delta_{0}\in(0,1)$ denotes the constant specified in Lemma˜2.3.
@@
$(\Delta+q_{0})(u_{f}-\tilde{u}_{f})=-\left(\tilde{q}(x,\lvert u_{f}\rvert)u_{f}-\tilde{q}(x,\lvert v_{f}\rvert)v_{f}\right)\text{ in $\Omega$},\quad\left.(u_{f}-\tilde{u}_{f})\right|_{\partial\Omega}=0.$
Using (2.16), we then obtain that
@@
$\quad\leq C(1+\lVert q_{0}\rVert_{L^{\infty}(\Omega)})\left(\begin{aligned} &\lVert\tilde{q}(\cdot,\lvert u_{f}\rvert)u_{f}-\tilde{q}(\cdot,\lvert v_{f}\rvert)v_{f}-q_{1}(\cdot)(\lvert u_{f}\rvert^{\alpha_{1}}u_{f}-\lvert v_{f}\rvert^{\alpha_{1}}v_{f})\rVert\\ &+\lVert q_{1}(\cdot)(\lvert u_{f}\rvert^{\alpha_{1}}u_{f}-\lvert v_{f}\rvert^{\alpha_{1}}v_{f})\rVert_{L^{\infty}(\Omega)}\end{aligned}\right).$
From (2.17b) and (2.18), the smallness of $\delta_{0}\in(0,1)$ ensures that
$\lVert u_{f}\rVert_{L^{\infty}(\Omega)}\leq 1\quad\text{and}\quad\lVert v_{f}\rVert_{L^{\infty}(\Omega)}\leq 1$
which allows us to apply (1.1b). Consequently, we use Lemma˜2.2 and the stability estimates in (2.17) and (2.18), to see that
@@ 3.2. An indicator functional and its approximation
We define $q_{b}(x,\lvert z\rvert):=q_{0}(x)+q_{1,b}\lvert z\rvert^{\alpha_{1}}$ for all $x\in\Omega$ and $z\in\mathbb{C}$, where $q_{1,b}\in L^{\infty}(\Omega)$ is the function given in (1.5a). By using Lemma˜2.3, there exists a $\delta_{0}\in(0,1)$ such that the following statement holds true: if $f\in U_{\delta_{0}}$, then there exists a unique solution $u_{b,f}\in\mathcal{A}^{2}(\Omega)$ to
@@
Using Lemma˜3.1, we know that there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1,b}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
@@
where $v_{f}$ is the unique solution to (2.1). Based on the indicator functional in [Kua24] and the structure of our current model, we propose the following variant: Ikehata’s enclosure method [Ike99, Ike00] suggests us to consider the following indicator functional:
$\mathcal{I}(f):=\int_{\partial\Omega}(\partial_{\nu}u_{f}-\partial_{\nu}\tilde{u}_{b,f})\overline{f}\,\mathrm{d}S\quad\text{for all $f\in U_{\delta_{0}}$},$
and the previous paper [Kua24] suggests us to define an auxiliary approximate indicator functional:
$\widetilde{\mathcal{I}}(f):=\int_{\partial\Omega}(\partial_{\nu}\tilde{u}_{f}-\partial_{\nu}\tilde{u}_{b,f})\overline{f}\,\mathrm{d}S\quad\text{for all $f\in U_{\delta_{0}}$.}$
@@
Let $\Omega$ be a bounded $C^{1,1}$-domain, and let $q$ satisfy (1.1b). Then there exists a constant $C=C(\Omega,\lVert q_{0}\rVert_{L^{\infty}(\Omega)},\lVert q_{1}\rVert_{L^{\infty}(\Omega)},\alpha_{1},C_{*})>0$ such that
@@
where $\delta_{0}\in(0,1)$ is the constant given in Lemma˜2.3.
Proof. Using the trace theorem [AH09, Theorem 7.3.11] (which requires condition $\partial\Omega\in C^{1,1}$), for each $f\in U_{\delta_{0}}$, we compute that
@@
Our lemma follows directly from Lemma˜3.1. ∎
@@
$\widetilde{\mathcal{I}}(f)=-\int_{D}q_{1,D}(x)\lvert v_{f}\rvert^{2+\alpha_{1}}\,\mathrm{d}x-\int_{\Omega}R(x,\lvert v_{f}\rvert)\lvert v_{f}\rvert^{2}\,\mathrm{d}x\quad\text{for all $f\in U_{\delta_{0}}$,}$
where $v_{f}$ is the unique solution to (2.1) and $\delta_{0}\in(0,1)$ denotes the constant specified in Lemma˜2.3.
@@
Proof of Lemma˜3.3. We see that the function $z_{f}:=\tilde{u}_{f}-\tilde{u}_{b,f}$ satisfies